OpenStudy (agent47):

Need to brush up on calc 2, have a few questions that I don't remember how to do. Please oblige :)

OpenStudy (agent47):

If C is a curve with parametric equations: x=2-e^t y=t^3+2t z=t -1<=t<=2, find parametric equations of the tangent line to C at t=1

OpenStudy (agent47):

@KingGeorge

OpenStudy (kinggeorge):

Let me review up on my parametric curves real quick.

OpenStudy (valpey):

The equation of a line in parametric form is: \[ Y = y(1)+y'(1)t\] \[ X = x(1)+x'(1)t\] \[ Z = z(1)+z'(1)t\]

OpenStudy (valpey):

But the parameter is offset, so for the tangent, t=0 is where it intersects the curve. I should have used a different parameter.

OpenStudy (valpey):

and \[y^{\prime}=\frac{dy}{dt}=3t^2+2\ \ etc.\]

OpenStudy (kinggeorge):

That \(z=t\) is throwing me off. I'm used to graphs in 2D.

OpenStudy (agent47):

But the parameter is offset, so for the tangent, t=0 is where it intersects the curve. I should have used a different parameter. what do you mean by this?

OpenStudy (valpey):

So if t is time, then the original curve starts at time = -1 below the xy plane and moves up as it curves around. At t=1 we take a snapshot of the point on the curve we are at and draw a tangent line in 3D. That tangent line can also be described parametrically. If I use the parameter s, say, we could describe a line that intersects where s=0 and the same time and place where t=1 on the first curve.

OpenStudy (valpey):

By using t as my parameter for the second curve, I give the impression that both the curve and the tanget line are being drawn at the same time, so to speak.

OpenStudy (agent47):

but i still just do X=x(1)+x'(1)t, etc. right?

OpenStudy (valpey):

Right, I just would have rather said X=x(1)+x'(1)s, etc. to emphasize that we are using a different parameter.

OpenStudy (valpey):

Alternatively, we could have done it like: X=x(1)-x'(1)+x'(1)t etc. and the curve and the tangent would have coincided at the same time and place.

OpenStudy (agent47):

okk i see, thanks a lot.